Sales and cost data for Van Buren Company. The total sales and cost information is based on the sale of 20,000 units.

Total Per Unit

Sales $570,000  $28.50

Variable costs 387,600 19.38

Fixed costs 140,000

Compute the total contribution margin, contribution margin ratio, and unit contribution margin for this company. Also, compute the increase in net income that will result from a $50,000 increase in sales and a 1,000-unit increase in sales.

1. Total contribution margin:

2. Contribution margin ratio:

3. Unit contribution margin:

4. Increase in net income from $50,000 increase in sales:

5. Increase in net income from 1,000-unit increase in sales:
Answer #1

Answers

--Requirement 1

A Sales $570,000
B Variable Costs $387,600
C = A - B Total Contribution margin $182,400

--Requirement 2

A Total Contribution margin $182,400
B Sales $570,000
C = (A/B) x 100 Contribution margin ratio 32%

--Requirement 3

A Total Contribution margin $182,400
B Total Units $20,000
C = A/B Unit Contribution margin $                     9.12

--Requirement 4

A Increase in Sales $ $50,000
B Contribution margin ratio 32%
C = A x B Increase in Net Income $16,000

--Requirement 5

A Increase in Unit Sales 1000
B Unit Contribution margin $                     9.12
C = A x B Increase in Net Income $9,120